Posting journal entry
160 Chapter 6 Accounting for Merchandising Businesses Illustrative Problem The following transactions were completed by Montrose Company during May of the current year. Montrose Compan y uses a perpetual inventory system. Purchased merchandise on account from Floyd Co., $4,000, terms FOB ship- ping point, 2/10, n/30, with prepaid freight of $120 added to the invoice. Purchased merchandise on account from Kramer Co., $8,500, terms FOB May 3. May 5. destination, 1/10, n/30. 6. Sold merchandise on account to C.F.Howell Co., list price $4,000, trade dis- count 30%, terms 2/10, n/30. The cost of the merchandise sold was si,125. 8. Purchased office supplies for cash, $150 10. Returned merchandise purchased on May 5 from Kramer Co., $1,300. 13. Paid Floyd Co. on account for purchase of May 3, less discount. 14. 15. Purchased merchandise for cash, $10,500. Paid Kramer Co. on account for purchase of May 5, less return of May 10 16. Received cash on account from sale of May 6 to C. F. Howell Co., less 19. Sold merchandise on MasterCard credit cards, $2,450. The cost of the mer- 22. Sold merchandise on account to Comer Co., $3,480, terms 2/10, n/30. The 24. Sold merchandise for cash, $4,350. The cost of the merchandise sold was 25. Received merchandise returned by Comer Co. from sale on May 22, S1A80 31. Paid a service processing fee of $140 for MasterCard sales and discount. discount. chandise sold was $980 cost of the merchandise sold was $1,400 $1,750. The cost of the returned merchandise was $600. 2. 1. If
Answer #1

SOLUTION:

Debit Credit
May-03 Merchandise Inventory 4,120
Accounts Payable–Floyd Co. 4,120
May-05 Merchandise Inventory 8,500
Accounts Payable–Kramer Co. 8,500
May-06 Accounts Receivable–C. F. Howell Co. 2,800
Sales 2,800
($4,000 - (30% * $4,000) = $2,800
May-06 Cost of Merchandise Sold 1,125
Merchandise Inventory 1,125
May-08 Office Supplies 150
Cash 150
May-10 Accounts Payable–Kramer Co. 1,300
Merchandise Inventory 1,300
May-13 Accounts Payable–Floyd Co. 4,120
Merchandise Inventory 40
Cash 4,040
(2% x $4,000 = $80; $4,000 - $80+$120 = $4,040
May-14 Merchandise Inventory 10,500
Cash
May-15 Accounts Payable–Kramer Co. 7,200
Merchandise Inventory 72
Cash 7,128
(8,500 -1,300) x 1% = $72; $8,500 - $1,300 - $72 = $7,128
May-16 Cash 2,744
Sales Discounts 56
Accounts Receivable–C. F. Howell Co. 2,800
$2,800 * 2% = $56; $2,800 - $56 = $2,744
May-19 Accounts Receivable–American Express 2,450
Sales 2,450
May-19 Cost of Merchandise Sold 980
Merchandise Inventory 980
May-22 Accounts Receivable–Comer Co. 3,480
Sales 3,480
May-22 Cost of Merchandise Sold 1,400
Merchandise Inventory 1,400
May-24 Cash 4,350
Sales 4,350
May-25 Sales Returns and Allowances 1,480
Accounts Receivable–Comer Co. 1,480
May-25 Merchandise Inventory 600
Cost of Merchandise Sold 600
May-31 Cash 2,310
Credit Card Expense 140
Accounts Receivable–American Express 2,450